OpenStudy (anonymous):

Which literal equations are equivalent to K - F/x ? Choose all answers that are correct. A. f=k x B. x= k f C. f= k/x D. x = f/k

OpenStudy (anonymous):

@ccswims

OpenStudy (ccswims):

For A and B are those still dividing or are you multiplying them?

OpenStudy (anonymous):

I don't get this at all.

OpenStudy (ccswims):

For A and B or you saying k*x and k*f or is it still k/x and k/f

OpenStudy (anonymous):

It doesn't change at all...It is still k/x and k/f...

OpenStudy (ccswims):

I believe everything except D

OpenStudy (anonymous):

Can you help me with another one?

OpenStudy (ccswims):

Yes

OpenStudy (anonymous):

Which literal equation is equivalent to d = 0.2t? A. t = D divided by 0.2 B. t = 0.2 divided by D C. 0.2 = t/d D. 0.2 = dt

OpenStudy (ccswims):

A

OpenStudy (anonymous):

Thanks! one more question please?

OpenStudy (ccswims):

sure!

OpenStudy (anonymous):

Solve for m. A. m = e+r/c B. m = er/c C. m = e - rc m = E – rc D. m = E + rc

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Um, what's the equation?

OpenStudy (ccswims):

Are there any equations with "m" not by itself before the answers?

OpenStudy (ccswims):

The final equation will have "m" by itself, but there should be an equation before with m by another variable

OpenStudy (igreen):

Can you post the question again? It looks scrambled.

OpenStudy (anonymous):

...

OpenStudy (igreen):

It looks like you didn't post the equation that we have to solve for..I just see options. @OpiGeode